OpenStudy (anonymous):

Given f(x) = –3x – 7, what is f(–4)? A. –1 B. –14 C. 5 D. 40

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm f(x) = –3x – 7 }\) ` E X A M P L E ` \(\large\color{midnightblue}{ \rm f(3) = –3(3) – 7 =-9-7=-16 }\) So, \(\large\color{midnightblue}{ \rm f(-4) = –3(-4)~ – 7 }\) ` S I M P L I F Y ` the given expression.